The length in centimeters of the 2 sides of a rectangle are consecutive odd intergers. The perimeter is 120 centimeters. whet is the area of the rectangle?
Answer by shweta(56)

Consecutive odd numbers are expressed as 2x+1, 2x+3
Perimeter of rectangle =120 cm
2(l+b)=120
2(2x+1+2x+3)=120
4x+4=120/2
4x+4=60
4x=60-4
4x=56
x=56/4
x=14
Two numbers are
2x+1= 2*14 +1= 28+1= 29
2x+3= 2*14 +3= 28+3= 31
Area of rectangle =l*b
=31*29
=899cm^2